OpenStudy (anonymous):

enter the solution of the system of equations below as an ordered pair. y+x=10 10y+20x=110

OpenStudy (ajprincess):

Multiply the first eqn by 10 then subtract t frm the sec eqn. Then u wil get an eqn in terms of x. solve for x. Substitute the value of x in the first eqn to find y.

OpenStudy (ajprincess):

can u do nw?

OpenStudy (anonymous):

should i multiply by -10?

OpenStudy (anonymous):

friend your answer is (1, 9) x=1 and y=9. please don't forget to click my answer as best if i am right.

OpenStudy (shubhamsrg):

lol @muhammad9t5 :D

OpenStudy (anonymous):

@shubhamsrg whats up bro?

OpenStudy (shubhamsrg):

sky !! :D

OpenStudy (shane_b):

Alternatively, you can do it this way: \[y+x=10\]\[10y+20x=110\]Put one of the variables in terms of the other. I'll use the first equation to do that:\[y=10-x\]Now substitute that equation into the second for every place you see a y:\[10(10-x)+20x=110\]Now solve for x. Once you have a value for x plug it back into the first equation to get y.

OpenStudy (shubhamsrg):

sorry,,lame pj,, :P we may chat through private messages buddy,,some people over here may get uselss notifications buddy,,which might irritate then after some time ;)

OpenStudy (anonymous):

you were right muhammad9t5.thank youu!

OpenStudy (shubhamsrg):

congo buddy,,you finally got a medal :) :P

OpenStudy (anonymous):

@Shane_B i also did like yours!

OpenStudy (shane_b):

Thanks. You got the answer of course but you should try to explain how you get them so people at least learn something :)